# Knowledge-guided machine learning for county-level corn yield prediction under drought

### v2

@@ 3.5.2 Drought-aware yield prediction loss function
Soil moisture is widely recognized to affect crop yield (Ines et al., 2013; Bushong et al., 2016) and numerous studies explored this relationship (Mladenova et al., 2017; Vergopolan et al., 2021; Pignotti et al., 2023). However most prior studies lacked an elegant quantitative approach addressing this issue. This study introduces a loss function adjusting predicted corn yield based on varying soil moisture levels.
The final yield prediction loss function was designed to improve the model’s accuracy while incorporating drought sensitivity and penalizing overestimation. It was formulated as:
@@
where $\bar{s}_{i}$ is the average soil moisture over the growing season for sample $i$, and $\varepsilon$ is a small constant to prevent numerical instability. We set $\varepsilon=1$ in our experiment. Since soil moisture plays a critical role in crop growth and yield formation, the loss function assigned a higher penalty to drier conditions, encouraging the model to be more responsive to soil moisture variations.
@@
This asymmetry discourages overestimation, particularly under drought conditions, where yield predictions tend to be more uncertain. By applying a stronger penalty to overestimated yields, the model is encouraged to be more conservative, reducing the risk of unrealistic predictions. We set $\lambda=2$ in our experiment.
@@ 3.5.3 Finetuning and testing with GEE county-level dataset
Following the pretraining of the model on the APSIM field-level dataset $\mathcal{D}_{\text{field/pretrain}}$, finetuning using the GEE county-level dataset $\mathcal{D}_{\text{county/finetune}}$ was required to improve the model for the county-level corn yield prediction task. This process involved initially partitioning the data into training, validation, and test sets, followed by predicting corn yield for the target years. Specific details regarding the partition and utilization of these datasets are elaborated upon in the experimental setup (Sec. 4.1). The loss functions $\mathcal{L}_{\text{yield}}$ and $\mathcal{L}_{\text{SM}}$ were also optimized on the training and validation datasets.
4 Experimental Results
@@ 4.1 Experimental setup
We implemented the DL models using the PyTorch framework (Paszke et al., 2019) and the traditional ML code with sklearn (Pedregosa et al., 2011). The models were run on A100-SXM4-40GB and A100-SXM4-80GB GPUs. For training, we used the ADAM optimizer (Kingma and Ba, 2014) with an initial learning rate of 0.001 and the ReduceLROnPlateau scheduler.
@@ 4.2.1 Specifying drought area
We investigated drought conditions in the United States (Drought.gov, 2024) during the corn growth season (June-September) (USDA, 2025) from 2019 to 2023 and present the corresponding maps (Fig. 3). In 2019 (Fig. 3(a)), drought conditions were predominantly observed in Iowa and Michigan. By 2020 (Fig. 3(b)), Iowa experienced severe drought, while moderate conditions affected North Dakota, South Dakota, Nebraska, and Illinois. Conditions deteriorated further in 2021 (Fig. 3(c)), with severe drought covering large parts of North Dakota, South Dakota, and Minnesota, while Nebraska and Iowa experienced moderate drought. In 2022 (Fig. 3(d)), severe drought persisted in Nebraska and Kansas, with South Dakota and Iowa experiencing moderate drought. In 2023 (Fig. 3(e)), severe drought was widespread across Nebraska, Kansas, Minnesota, Iowa, and Wisconsin, while Missouri faced moderate drought conditions.
@@ 4.2.3 Drought impact on corn yield
Finally, we examined the specific impact of drought on corn yield. We plotted the difference in corn yield for each year from 2019 to 2023 relative to the previous year’s yield (Fig. 5), where negative values indicate a yield reduction. In 2019 (Fig. 5(a)), drought-affected areas included Iowa and Michigan; however, significant yield reductions were observed only in Michigan. In 2020 (Fig. 5(b)), Iowa faced drought along with a derecho storm (Hosseini et al., 2020), leading to noticeable yield losses. In 2021 (Fig. 5(c)), drought impacted North Dakota, South Dakota, Minnesota, Nebraska, and Iowa, yet significant yield reductions occurred only in North Dakota and South Dakota. In 2022 (Fig. 5(d)), drought-affected states included Nebraska, Kansas, South Dakota, and Iowa, but yield declines were primarily observed in Nebraska, Kansas, and South Dakota. In 2023 (Fig. 5(e)), although Nebraska and Kansas continued to experience drought, corn yield did not decrease significantly compared to the previous year, likely due to the already reduced yield in 2022. However, Iowa and Missouri experienced some yield losses due to ongoing drought conditions. Thus, we concluded that drought consistently led to a decline in corn yield.
@@ 4.2.4 Statistical summary of drought-affected states
Based on the above statistical analysis, we summarized the drought-affected areas that experienced yield reductions in Table 7. In the experiment, we focused on these states that were affected by drought and experienced yield reductions.
@@
Year Area type Drought-affected states Obviously yield-reduced states
2019 IA,MI MI
2020 IA,ND,SD,NE,IL IA
2021 ND,SD,MN,NE,IA ND,SD
2022 NE,KS,SD,IA NE,KS,SD
2023 NE,KS,MN,IA,WI,MO IA,MO
@@ 4.3 Evaluation results
Year Method KGML-SM LR MLP RR RF
@@ 4.4 Ablation study of different components in KGML-SM model
To further demonstrate the contribution of each module in our KGML-SM model, we conducted a series of ablation studies. We began with a baseline model trained using only the GEE county-level dataset and the attention module (Att) (Wang et al., 2025). Next, we incorporated the APSIM field-level dataset pretraining to evaluate its impact on model performance (Att+sim). We then introduced the W2S encoder to assess its effectiveness in integrating soil moisture information (Att+sim+W2S). Finally, the impact of the drought-aware loss function was evaluated by comparing the performance of the KGML-SM model with the (Att+sim+W2S) model. We specifically compared the differences among models and their components in Table 9.
@@
Method Attention module Field-level data pretraining W2S encoder Drought-aware loss
@@
Through the ablation study of all model components (Table 10), we found that the APSIM field-level dataset pretraining contributed the most to performance improvement. This indicated that our APSIM field-level dataset effectively captured county-level data patterns, playing a crucial role in enabling the model to learn the relationship between corn yield and agricultural variables. Additionally, our drought-aware loss function significantly enhanced performance in most years, demonstrating that incorporating prior knowledge of the relationship between drought and corn yield can effectively guide the training and prediction of ML models.
@@
Table 10: Ablation study of different components in KGML-SM model
Year Method KGML-SM Att+sim+W2S Att+sim Att
RMSE R2 RMSE R2 RMSE R2 RMSE R2
2019 0.964 0.741 0.974 0.732 1.087 0.715 1.268 0.57
2020 0.98 0.792 0.981 0.783 1.003 0.77 1.011 0.77
2021 1.104 0.836 1.127 0.811 1.143 0.814 1.195 0.808
2022 1.085 0.837 1.119 0.802 1.144 0.802 1.315 0.779
2023 1.071 0.807 1.101 0.805 1.114 0.781 1.201 0.766
4.5 Spatial Visualization of model performance
In Fig. 6, we plotted the error maps of the prediction results from 2019 to 2023 for the two best performed models: KGML-SM and RF model. In 2019 (Fig. 6(a)), the lower part of Michigan showed noticeable overestimation in the RF model, whereas our KGML-SM model performed well. In 2020 (Fig. 6(b)), although the derecho storm (Hosseini et al., 2020) in Iowa also had some impact on our KGML-SM model, we found that the RF model exhibited more severe overestimation. In 2021 (Fig. 6(c)), overestimation in the upper-left region, particularly in North Dakota and South Dakota, was significantly reduced. In 2022 (Fig. 6(d)), overestimation decreases notably in North Dakota, South Dakota, and Kansas, though the effect in Kansas was less pronounced. In 2023 (Fig. 6(e)), Iowa showed a reduction in overestimation, while the effect in Missouri was less noticeable. Overall, our model shows a notable improvement in addressing the issue of overestimation in drought-affected areas.
@@ 4.5 Spatial Visualization of model performance
4.6 Understanding model bias and prediction-observation alignment
To analyze model prediction performance, we generated scatter plots for the KGML-SM model and the RF model (Fig. 7). These plots helped visualize the relationship between observed and predicted values, revealing patterns of overestimation, underestimation, and potential prediction biases across different years. In 2019 (Fig. 7(a)), the KGML-SM model presented a noticeably narrower distribution, indicating a lower spread in prediction errors. In 2020 (Fig. 7(b)), the predictions of the KGML-SM model were noticeably more concentrated along the diagonal and exhibited symmetry on both sides, whereas the RF model produced more dispersed predictions in high-yield regions. In 2021 (Fig. 7(c)), the RF model exhibited prediction collapse, where certain observed values correspond to nearly identical predicted values, likely due to overfitting or insufficient variability in learned representations. In 2022 (Fig. 7(d)) and 2023 (Fig. 7(e)), the KGML-SM model maintained a narrower and more concentrated prediction distribution. This comparison highlighted the advantage of the KGML-SM model in mitigating prediction collapse and improving overall robustness across different years.
Figure 7: The scatter plot for (1) KGML-SM and (2) RF model from (a) 2019 to (e) 2023.

### v3

@@ 3.5.2 Drought-aware yield prediction loss function
Soil moisture is widely recognized to affect crop yield (Ines et al., 2013b; Bushong et al., 2016) and numerous studies explored this relationship (Mladenova et al., 2017; Vergopolan et al., 2021; Pignotti et al., 2023). However most prior studies lacked an elegant quantitative approach addressing this issue. This study introduces a loss function adjusting predicted corn yield based on varying soil moisture levels. The final objective function jointly optimizes both components and is defined as:
$\mathcal{L}_{total}=\mathcal{L}_{SM}+\mathcal{L}_{yield}$ (8)
The yield prediction loss function was designed to improve the model’s accuracy while incorporating drought sensitivity and penalizing overestimation. It was formulated as:
@@
where $\bar{s}_{i}$ is the average soil moisture during the growing season (April–October) for sample $i$, calculated from the input soil moisture dataset (APSIM-simulated data for pretraining, and GEE/SMAP soil moisture for finetuning). $\varepsilon$ is a small constant to prevent numerical instability, which we set to 1 in our experiment. Since soil moisture plays a critical role in crop growth and yield formation, the loss function assigned a higher penalty to drier conditions, encouraging the model to be more responsive to soil moisture variations.
@@
This asymmetry discourages overestimation, particularly under drought conditions, where yield predictions tend to be more uncertain. By applying a stronger penalty to overestimated yields, the model is encouraged to be more conservative, reducing the risk of unrealistic predictions. The value of hyperparameter $\lambda$ was selected based on validation performance, and we set $\lambda=2$ in our experiment (Sec. 4.2).
@@ 3.5.3 Finetuning and testing with GEE county-level dataset
Following the pretraining of the model on the APSIM field-level dataset $\mathcal{D}_{\text{field/pretrain}}$, finetuning using the GEE county-level dataset $\mathcal{D}_{\text{county/finetune}}$ was required to improve the model for the county-level corn yield prediction task. This process involved initially partitioning the data into training, validation, and test sets, followed by predicting corn yield for the target years. We ensured independent validation by adopting a temporal split strategy, where models were trained on preceding years and tested on the target year. Specific details regarding the partition and utilization of these datasets are elaborated upon in the experimental setup (Sec. 3.6). The loss functions $\mathcal{L}_{\text{yield}}$ and $\mathcal{L}_{\text{SM}}$ were also optimized on the training and validation datasets.
@@ 3.6 Experimental setup
We implemented the DL models using the PyTorch framework (Paszke et al., 2019) and the traditional ML code with sklearn (Pedregosa et al., 2011). The models were run on A100-SXM4-40GB and A100-SXM4-80GB GPUs. For pretraining, we used a batch size of 64, a learning rate of 0.001, the Adam optimizer (Kingma and Ba, 2014), and the ReduceLROnPlateau scheduler with a patience of 5, with training stopped once the RMSE dropped below 1. For finetuning, we used a smaller batch size of 16 but kept the same learning rate, optimizer, and scheduler, and applied early stopping based on validation loss. The full configurations are summarized in Table 6.
Table 6: Training configurations for pretraining and finetuning.
Stage
Batch
size
Learning
rate
Optimizer Scheduler Epochs Loss function Stopping criterion
Pretrain 64 0.001 Adam
ReduceLROnPlateau
(patience=5)
$\sim$50 MSE Training RMSE << 1
Finetune 16 0.001 Adam
ReduceLROnPlateau
(patience=5)
$\sim$30 MSE Early stopping based on validation loss
@@ 3.7.1 Specifying drought area
To objectively determine drought conditions, we used county‐level data from the U.S. Drought Monitor (USDM) (Drought.gov, 2024). For each Corn Belt state during the corn growing season (June–September; (USDA, 2025)), the USDM provides categorical drought classifications ranging from None to D4 (Exceptional Drought). Fig. 3 shows the spatial distribution of these drought categories, providing a visual representation of drought intensity across years. To quantitatively assess drought severity, we further aggregated the USDM records of county areas under each drought category (None–D4). By summing the areas across drought categories, we calculated the proportion of each state’s total area that experienced drought in a given year. To complement the maps, Table 7 reports the state‐level proportions of drought‐affected counties, with values exceeding 30% highlighted in bold to emphasize years and regions with particularly widespread drought.
@@
Table 7: Proportion of counties affected by drought (categories D0–D4) in each of the 12 U.S. Corn Belt states (2018–2023), derived from U.S. Drought Monitor (USDM) county-level records during the corn growing season (June–September). Values indicate the fraction of state area under drought, with those exceeding 30% highlighted in bold.
Year IA IL IN KS MI MN MO ND NE OH SD WI
2019 17% 16% 15% 9% 15% 11% 1% 37% 1% 4% 0% 1%
2020 64% 19% 49% 64% 22% 48% 19% 68% 55% 47% 59% 8%
2021 89% 25% 18% 30% 65% 100% 8% 100% 76% 12% 100% 66%
2022 62% 38% 47% 88% 30% 23% 61% 8% 97% 9% 73% 37%
2023 100% 94% 73% 97% 72% 96% 95% 57% 93% 58% 61% 98%
@@ 3.7.3 Drought impact on corn yield
Finally, we examined the interannual impact of drought on corn yield. Fig. 5 shows the county-level yield differences between each year from 2019 to 2023 and the previous year, where negative values indicate yield reductions. To further quantify the extent of yield losses, we calculated for each state the proportion of counties experiencing a yield decline greater than 1 t/ha relative to the previous year. These results are summarized in Table 8, with values above 30% highlighted in bold to emphasize states and years with widespread yield reductions.
@@
Table 8: Proportion of counties with yield reductions greater than 1 t/ha relative to the previous year across 12 Corn Belt states (2019–2023), with those exceeding 30% highlighted in bold.
Year IA IL IN KS MI MN MO ND NE OH SD WI
2019 27% 78% 67% 15% 21% 36% 10% 50% 32% 54% 42% 30%
2020 54% 1% 2% 20% 8% 5% 17% 33% 16% 7% 4% 7%
2021 1% 0% 4% 16% 0% 50% 42% 87% 3% 0% 71% 0%
2022 33% 1% 30% 76% 42% 3% 28% 0% 85% 26% 39% 12%
2023 21% 23% 1% 15% 11% 30% 36% 17% 8% 0% 2% 17%
@@ 3.7.4 Statistical summary of drought-affected states
Based on the above statistical analysis, we summarized the drought-affected areas that experienced yield reductions in Table 9. This table shows the states where drought and yield loss overlapped. In 2019, drought was limited, but many counties still had large yield reductions. This was mainly due to delayed planting and record prevented planting caused by heavy rainfall and flooding (American Farm Bureau Federation, 2019; Farm Service Agency, U.S. Department of Agriculture, 2019). In 2020, fewer yield losses were observed, likely because many states had already lost yield in 2019, leaving less room for further decline. In 2021 and 2022, drought-affected states also showed clear yield reductions, which matches the expected pattern. In 2023, drought was widespread across all states, but only a few states had obvious yield losses, because severe drought in 2022 had already caused large yield reductions.
@@
Year Area type Drought-affected states Obviously yield-reduced states Intersection
2019 ND IL,IN,MN,ND,NE,OH,SD,WI ND
2020 ND,IA,KS,SD,NE,MN,IN,OH IA,ND IA,ND
2021 MN,ND,SD,IA,NE,WI,MI MN,MO,ND,SD MN,ND,SD
2022 NE,KS,SD,IA,MO,IN,IL,WI IA,IN,KS,MI,NE,SD IA,IN,KS,NE,SD
2023 All the states MN,MO MN,MO
4 Experimental Results
@@ 4.1 Evaluation results
Year Method KGML-SM LR MLP $R$ RF
@@ 4.2 Ablation study of different components in KGML-SM model
To further demonstrate the contribution of each module in our KGML-SM model, we conducted a series of ablation studies. We began with an attention-based baseline model without soil moisture inputs (Att w/o SM), which used all features except soil moisture. We then included soil moisture in the input features to form the Att model. Next, we incorporated the APSIM field-level dataset for pretraining (Att+sim), followed by the addition of the W2S encoder to integrate soil moisture dynamics (Att+sim+W2S). To further analyze the contribution of different loss components in our drought-aware loss $\mathcal{L}_{\text{yield}}=\frac{1}{N}\sum_{i=1}^{N}d_{i}\left[(y_{i}-\hat{y}_{i})^{2}+\lambda\max(0,\hat{y}_{i}-y_{i})^{2}\right]$, we first added the Soil Moisture Weighted (SMW) term, represented by $d_{i}$, which assigns greater weight to errors under low soil moisture conditions (Att+sim+W2S+SMW). We then introduced the Overestimation penalty (OE) term, $\lambda\max(0,\hat{y}_{i}-y_{i})^{2}$, which penalizes yield overestimation more strongly, particularly under drought conditions. Combining both SMW and OE yielded the full KGML-SM framework. A summary of the components of each ablation model is provided in Table 11.
@@
Method
Attention
module
Soil
moisture
Field-level data
pretraining
W2S
encoder
SMW
loss
OE
loss
Att w/o SM
@@
Through the ablation study of all model components (Table 12), we found that the APSIM field-level dataset pretraining contributed the most to performance improvement. This indicated that our APSIM field-level dataset effectively captured county-level data patterns, playing a crucial role in enabling the model to learn the relationship between corn yield and agricultural variables. When comparing the baseline attention model without soil moisture (Att w/o SM) to the one including soil moisture inputs (Att), the performance difference was marginal across most years. This result suggests that simply adding raw soil moisture values provides limited benefits, highlighting the necessity of our subsequent modules to better exploit soil moisture information.
Additionally, our drought-aware components also contributed to performance gains. While the SMW Loss improved performance in some drought-affected years, the effect was not consistent. This is because overestimation of yield can arise not only from drought but also from other natural hazards such as flooding, extreme temperatures, or pest outbreaks, which are not fully captured by soil moisture weighting alone. In contrast, the overestimation penalty directly constrains the model against overprediction across diverse adverse conditions, thereby providing more robust improvements.
@@
Table 12: Ablation study of different components in KGML-SM model (RMSE and R2 across years).
Method Year 2019 2020 2021 2022 2023
RMSE R2 RMSE R2 RMSE R2 RMSE R2 RMSE R2
Att w/o SM 1.258 0.578 1.053 0.752 1.149 0.812 1.315 0.780 1.096 0.784
Att 1.268 0.570 1.011 0.770 1.195 0.808 1.315 0.779 1.201 0.766
Att+sim 1.087 0.715 1.003 0.770 1.143 0.814 1.144 0.802 1.114 0.781
Att+sim+W2S 0.974 0.732 0.981 0.783 1.127 0.811 1.119 0.802 1.101 0.805
Att+sim+W2S+SMW 1.097 0.711 1.054 0.763 1.112 0.832 1.074 0.821 1.043 0.812
KGML-SM 0.964 0.741 0.980 0.792 1.104 0.836 1.085 0.837 1.071 0.807
In DL, loss functions often include hyperparameters that control the relative importance of different error components (Goodfellow et al., 2016). In the drought-aware loss, the coefficient $\lambda$ determines the penalty strength applied to yield overestimation. Choosing an appropriate value of $\lambda$ is therefore critical: too small a value would fail to constrain overestimation effectively, whereas too large a value could distort the optimization and harm overall accuracy. To ensure a principled selection, we performed hyperparameter tuning on the validation set by testing multiple candidate values ($\lambda=0,1,2,5,10$) rather than assigning it arbitrarily. The results are summarized in Table 13. We found that $\lambda=2$ consistently offered the best trade-off between reducing RMSE and improving $R^{2}$, while both smaller and larger values led to inferior performance. Consequently, we adopted $\lambda=2$ as the default setting in the KGML-SM framework.
Table 13: Ablation study on different values of the overestimation penalty coefficient $\lambda$ (RMSE and R2 for 2019–2023). The best $\lambda$ was selected based on validation performance.
Year $\lambda=0$ $\lambda=1$ $\lambda=2$ $\lambda=5$ $\lambda=10$
RMSE R2 RMSE R2 RMSE R2 RMSE R2 RMSE R2
2019 1.097 0.711 0.982 0.752 0.964 0.741 1.032 0.730 1.115 0.695
2020 1.054 0.763 0.976 0.775 0.980 0.792 1.021 0.749 1.109 0.710
2021 1.112 0.832 1.122 0.819 1.104 0.836 1.166 0.785 1.235 0.745
2022 1.074 0.821 1.108 0.829 1.085 0.837 1.153 0.792 1.223 0.755
2023 1.043 0.812 1.062 0.801 1.071 0.807 1.184 0.776 1.246 0.740
To evaluate the effect of filtering the simulated dataset (Sec. 3.3.2), we compared KGML-SM trained with and without dataset filtering. As shown in Table 14, filtering substantially improved model performance across multiple years. This demonstrates that eliminating unreliable simulation samples is critical for constructing a robust pretraining dataset and enhancing generalization to county-level observations.
Table 14: Effect of filtering the simulated dataset on model performance (RMSE and R2 for 2019–2023).
Method 2019 2020 2021 2022 2023
RMSE R2 RMSE R2 RMSE R2 RMSE R2 RMSE R2
Unfiltered 1.120 0.705 1.010 0.768 1.198 0.807 1.305 0.775 1.225 0.760
Filtered (KGML-SM) 0.964 0.741 0.980 0.792 1.104 0.836 1.085 0.837 1.071 0.807
4.3 Prediction error spatialization and model bias
In Fig. 6, we present error maps of prediction results from 2019 to 2023 for the two best-performing models: KGML-SM and the RF baseline. Across all five years, KGML-SM consistently reduced overestimation compared with RF. In 2019, while overestimation in drought-affected areas of North Dakota was not substantially alleviated, several eastern states showed reductions. This is consistent with the fact that yield losses in 2019 were mainly due to delayed planting caused by prevented planting (American Farm Bureau Federation, 2019; Farm Service Agency, U.S. Department of Agriculture, 2019), rather than drought. In 2020, although the derecho storm (Hosseini et al., 2020) in Iowa also posed challenges for KGML-SM, the RF model exhibited much more severe overestimation, indicating that KGML-SM is effective not only under drought conditions but also in mitigating overestimation associated with other natural hazards. The improvements were particularly pronounced in 2021 and 2022, when RF showed widespread overestimation across the northwestern Corn Belt, corresponding to regions of drought-induced yield reduction in 2021, and across Kansas and Nebraska in 2022. In 2023, KGML-SM also reduced overestimation in Minnesota, further demonstrating its robustness. These results show the robustness of KGML-SM in mitigating systematic yield overestimation across diverse climatic conditions.
@@ 4.3 Prediction error spatialization and model bias
To analyze model prediction performance, we generated scatter plots for the KGML-SM model and the RF model (Fig. 7). These plots helped visualize the relationship between observed and predicted values, revealing patterns of overestimation, underestimation, and potential prediction biases across different years. In 2019 (Fig. 7(a)), the KGML-SM model presented a noticeably narrower distribution, indicating a lower spread in prediction errors. In 2020 (Fig. 7(b)), the predictions of the KGML-SM model were noticeably more concentrated along the diagonal and exhibited symmetry on both sides, whereas the RF model produced more dispersed predictions in high-yield regions. In 2021 (Fig. 7(c)), the RF model exhibited prediction collapse, where certain observed values corresponded to nearly identical predicted values, likely due to overfitting or insufficient variability in learned representations. In 2022 (Fig. 7(d)) and 2023 (Fig. 7(e)), the KGML-SM model maintained a narrower and more concentrated prediction distribution. This comparison highlighted the advantage of the KGML-SM model in mitigating prediction collapse and improving overall robustness across different years.
Figure 7: Scatter plots of predicted versus observed county-level corn yields for KGML-SM (top row) and the RF model (bottom row) from 2019 to 2023. Panels (a1)–(e1) correspond to KGML-SM predictions for 2019–2023, while panels (a2)–(e2) correspond to RF predictions for the same years. The horizontal axis represents the reported (observed) yield (t/ha), and the vertical axis represents the model-predicted yield (t/ha). The color scale indicates the density of counties, and the dashed line shows the 1:1 reference line for comparison.